OpenStudy (anonymous):

Find a cubic function with the given zeros. 6, -5, 2

OpenStudy (goformit100):

(-b/a)x^3 + (c/a)x^2 + (d/a)x + f

OpenStudy (anonymous):

Im still not sure how to do this

OpenStudy (anonymous):

Can you do it and ill see if i understand all the steps?

OpenStudy (goformit100):

\[\alpha x^3 +\beta x^2 + \gamma x +\delta =0\]

OpenStudy (anonymous):

how do i know what a b and y are?

hero (hero):

The given zeroes: x = 6 x = -5 x = 2 subtract each number from both sides: x - 6 = 0 x +5 = 0 x - 2 = 0 Therefore the factors are (x - 6), (x + 5) and (x - 2) Multiply them together to get the cubic

OpenStudy (anonymous):

wait how do i multiply them together?

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

is this the answer? f(x) = x3 + 3x2 - 28 + 60

hero (hero):

(x - 6)(x + 5) = x(x + 5) - 6(x + 5) = x^2 +5x - 6x - 30 = x^2 -x - 30 x^2 - x - 30(x - 2) = x^2(x - 2) - x(x - 2) - 30(x - 2) = x^3 - 2x^2 - x^2+ 2x - 30x + 60 = x^3 - 3x^2 - 28x + 60

OpenStudy (anonymous):

Yay! i did do it right